# DriVerse: Navigation World Model for Driving Simulation via Multimodal Trajectory Prompting and Motion Alignment

### v1

@@ 1 Introduction
Other approaches like GAIA-1 [29], MUVO [3], ADriver-I [33], DriveDreamer-2 [76], DriveWorld [45], BEVWorld [74], WorldGPT [16] and HERMES [78] rely on text prompts or discrete control signals. While effective for open-ended generation, they lack precise control over the generation process for specific trajectories, making them less suitable for evaluation or planning under strict constraints. Furthermore, trajectory-conditioned models such as Vista [14], OccWorld [77], OccSora [56], DOME [17], and Imagine-2-Drive [15] attempt to encode 3D trajectories as generation conditions. However, directly encoding 3D coordinates or projecting them onto 2D planes often results in poor alignment with pre-trained generative features. This misalignment makes it difficult to learn from limited data and can degrade the generalization capability of the underlying generative backbone. As a result, these methods struggle to produce temporally consistent videos in challenging scenarios such as turning or lane changing.
@@ 2.1 Video Generation Models
To effectively utilize powerful pre-trained image generators like Stable Diffusion [50], subsequent works have integrated temporal modules into the 2D backbone and fine-tuned them on large-scale video datasets [1]. Other notable directions in this field include scalable Transformer-based architectures [44], modeling in spatiotemporally compressed latent spaces (as seen in W.A.L.T.[21] and Sora[4]), and discrete token generation guided by language models [40, 6]. For a more detailed overview, we recommend readers consult [49]. Recently, models like Stable Video Diffusion (SVD)[2], based on U-Net, and Wan2.1[62], leveraging Diffusion Transformer (DiT) architecture for better scale-up performance, have further refined the pipeline with enhanced training methodologies and data selection strategies.
In the realm of controllable video generation, several studies have integrated additional control signals, such as depth maps and pose sequences, to precisely guide scene and human movements [20, 9, 75, 38, 31, 66]. For example, SparseCtrl [20] employs sketch images as spatial control inputs, resulting in enhanced video quality and improved temporal coherence. In terms of camera control, AnimateDiff utilizes lightweight LoRA-based fine-tuning [30] to accommodate various types of camera movements.
@@ 2.2 Navigation World Models for Driving
World models aim to predict future scene evolution based on current observations [22, 18, 54], and have increasingly become a research focus in autonomous driving. Many approaches [39, 43, 41, 63, 70] focus on predicting 2D visual representations, demonstrating strong capabilities in dynamic modeling.
@@
However, these world models often rely on high-level semantic reasoning and struggle to produce physically plausible, detail-rich long-horizon videos when ego-vehicle behaviors change drastically (e.g., turning or lane changing). Vista [14], built upon GenAD, incorporates extensive additional data and introduces structural and motion losses to improve dynamic modeling and structural fidelity, achieving high-resolution, high-quality long-term scene generation. Subsequent works further advance the duration of predictions [19, 32] and unify prediction with planning [11, 59] or perception [42].
@@ 3.1 Multimodal Trajectory Prompting
[…]
@@ 3.3 Dynamic Window-based Generation
DriveGAN [39] CVPR 21 nuScenes 256$\times$256 73.4 502.3
DriveDreamer [57] ECCV 24 nuScenes 128$\times$192 12Hz, 4s 52.6 452.0
WoVoGen [43] ECCV 24 nuScenes 256$\times$448 2Hz, 2.5s 27.6 417.7
GenAD [67] CVPR 24 nuScenes 256$\times$448 2Hz, 4s 15.4 184.0
Drive-WM [59] CVPR 24 nuScenes 192$\times$384 2Hz, 8s 15.2 122.7
Vista [14] NeurIPS 24 nuScenes 480$\times$832 10Hz, 15s 23.4 158.0
DriVerse (Ours) nuScenes 480$\times$832 10Hz, 8s 18.2 95.2
DriVerse (Ours) nuScenes 480$\times$832 10Hz, 15s 20.1 143.5
Vista NeurIPS 24 Waymo Open Dataset 480$\times$832 10Hz, 15s 29.2 285.7
DriVerse (Ours) Waymo Open Dataset 480$\times$832 10Hz, 8s 25.0 204.3
DriVerse (Ours) Waymo Open Dataset 480$\times$832 10Hz, 15s 28.4 264.5
DriVerse (Ours) Waymo Open Dataset 480$\times$832 10Hz, 30s 31.3 281.4

### v2

@@ 1 Introduction
Other approaches like GAIA-1 [31], MUVO [4], ADriver-I [36], DriveDreamer-2 [89], DriveWorld [52], BEVWorld [87], MPDrive [88], WorldGPT [18] and HERMES [91] rely on text prompts or discrete control signals. While effective for open-ended generation, they lack precise control over the generation process for specific trajectories, making them less suitable for evaluation or planning under strict constraints. Furthermore, trajectory-conditioned models such as Vista [16], OccWorld [90], OccSora [67], DOME [19], and Imagine-2-Drive [17] attempt to encode 3D trajectories as generation conditions. However, directly encoding 3D coordinates or projecting them onto 2D planes often results in poor alignment with pre-trained generative features. This misalignment makes it difficult to learn from limited data and can degrade the generalization capability of the underlying generative backbone. As a result, these methods struggle to produce temporally consistent videos in challenging scenarios such as turning or lane changing.
@@ 2.1 Video Generation Models
To effectively utilize powerful pre-trained image generators like Stable Diffusion [58], subsequent works have integrated temporal modules into the 2D backbone and fine-tuned them on large-scale video datasets [1]. Other notable directions in this field include scalable Transformer-based architectures [51], modeling in spatiotemporally compressed latent spaces (as seen in W.A.L.T.[23] and Sora[5]), and discrete token generation guided by language models [43, 7]. Effective factors in visual in-context learning have also been analyzed in detail [63]. For a more detailed overview, we recommend readers consult [56]. Recently, models like Stable Video Diffusion (SVD)[2], based on U-Net, and Wan2.1[73], leveraging Diffusion Transformer (DiT) architecture for better scale-up performance, have further refined the pipeline with enhanced training methodologies and data selection strategies.
In the realm of controllable video generation, several studies have integrated additional control signals, such as depth maps and pose sequences, to precisely guide scene and human movements [22, 10, 86, 41, 33, 78]. Complementary to diffusion video generators, autoregressive next-focus prediction over images [45] and the framing of image editing as a temporally degenerate process [44] define alternative inductive biases for structured appearance change, while 3D Gaussian score-distillation editing [57] and NeRF-based multi-view object detection indoors [35] stress cross-view geometry in settings complementary to monocular I2V on streets. Visual reference prompts for foundation segmenters (e.g., SAM [62]) further highlight how non-textual conditions steer large vision models. For example, SparseCtrl [22] employs sketch images as spatial control inputs, resulting in enhanced video quality and improved temporal coherence. In terms of camera control, AnimateDiff utilizes lightweight LoRA-based fine-tuning [32] to accommodate various types of camera movements.
@@ 2.2 Navigation World Models for Driving
World models aim to predict future scene evolution based on current observations [24, 20, 65], and have increasingly become a research focus in autonomous driving. Many approaches [42, 50, 47, 74, 81] focus on predicting 2D visual representations, demonstrating strong capabilities in dynamic modeling. Closely related, learning multiple probabilistic decisions from a latent world model [77] further links long-horizon rollouts to action diversity. In parallel, vision–language research revisits MLLM image classification under classic benchmarks [49], improves descriptive captioning for multimodal perception with specialist models [64], explores grow-and-refine memory for agentic multimodal learning [3], and mitigates visual forgetting in efficient MLLMs via resampling [14]—all adjacent trends for scalable conditioning interfaces in embodied and automotive AI, though with objectives beyond trajectory-consistent video.
@@
However, these world models often rely on high-level semantic reasoning and struggle to produce physically plausible, detail-rich long-horizon videos when ego-vehicle behaviors change drastically (e.g., turning or lane changing). Vista [16], built upon GenAD, incorporates extensive additional data and introduces structural and motion losses to improve dynamic modeling and structural fidelity, achieving high-resolution, high-quality long-term scene generation. Subsequent works further advance the duration of predictions [21, 34] and unify prediction with planning [12, 71] or perception and localization [48, 46].
@@ 3.1 Multimodal Trajectory Prompting
Trend Tokenization of 3D Trajectories. To encode directional priors of future trajectories, we discretize each 3D trajectory segment into a predefined trend vocabulary. Specifically, the $360^{\circ}$ horizontal plane is uniformly divided into 12 angular sectors, each corresponding to a directional token $\texttt{<T}_{k}\texttt{>}$, where $k\in\{1,\dots,12\}$ denotes a clock-hour direction. Given a trajectory $\{\mathbf{x}_{t}\}_{t=1}^{T}$, we compute segment directions $\Delta\mathbf{x}_{t}=\mathbf{x}_{t+1}-\mathbf{x}_{t}$ and map each to its token via the quantized azimuthal angle $\theta_{t}$:
@@ 3.3 Dynamic Window-based Generation
DriveGAN [42] CVPR 21 nuScenes $\times256$ 73.4 502.3
[…]